solve for theta if tan theta = 2 sin theta on 0 degrees < or = theta < or = 360 degrees.
Answer by Alan3354(31532)

tan = 2*sin
sin/cos = 2*sin
1/cos = 2
cos = 1/2
theta = 60º, 300º